calculate (to three significant figures) the numerical part of the conversion factors needed to carry out the following unit conversions:
a) Density in 1.00 g/cm^3 to kg/m^3
b) area in 1.00 km^2 to cm^2
c) mass/area of aluminum foil in 1.00 mg/cm^2 to g/m^2
Answer #1

a) 1.00 g/cm3 = 1.00 (10-3 kg) / (10-6 m3) = 1000 kg / m3

b) 1.00 km2 = 1.00 (1010 cm2) = 1.00 * 105 cm2

c) 1.00 mg/cm2 = 1.00 * 10-3 g / (10-4 m2) = 10.00 g/m2